OpenStudy (anonymous):

in parallelogram ABCD, angle A contains 60 degree. What is the sum of angle B and angle D? 1. 60 degree 2. 180 degree 3. 240 degree 4. 280 degree 5. 300 degree

OpenStudy (anonymous):

In a parallelogram, diagonal's are congruent. If thats the case, then if a = 60 then c is also 60 since it is a's diagonal. Both diagonals added together equal 120. Subtract the total of those two diagonals from the total of 360 360 - 120 = 240 240 is the sum of both B and D, the two remaining diagonals. B=120 D= 120

OpenStudy (anonymous):

Also, if you wanted extra info, Divide 240 by the two remaining sides, which are B & D 240/2 = 120 120 is the angle of the two remaining diagonals also

OpenStudy (anonymous):

how is mine wrong =p 120 +120 +60+60 =360 too >.>

OpenStudy (anonymous):

Nooo, if a is 60 right?? then its diagonal, c, is also 60. this together is 120 soo 360 -120 is 240 right. 240 is the some of diagonals B and D Divided its 120....

OpenStudy (anonymous):

meh, just nvm lol ima confuse my head when im to sleepy

OpenStudy (anonymous):

Or basically the asnwer is 3, which is what both of us got in the end...

OpenStudy (anonymous):

exr0sis is correct

OpenStudy (anonymous):

My bad silly mistake. exorsis is correct.